OpenStudy (anonymous):

6x - 2y = 12 the slope (m) of this equation is ___ ? *hint* you first need to put the equation in slope intercept form.

sam (.sam.):

3

sam (.sam.):

6x-2y=12 -2y=-6x+12 y=3x-6

OpenStudy (anonymous):

the slope is going to be the coefficent of x, in this case 3

OpenStudy (anonymous):

oh ok i see